Grade 9Algebra

find the relation between x and y in order that the rth Arithemetic Mean between x and 2y may be the same as the rth mean between 2x and y,n means being inserted in each case

PRAKHAR  GUPTA
12 Years agoGrade 9

1 Answer

Pratik Tibrewal
12 Years ago
rth arithmetic mean b/w x and 2y : x + r*((2y-x)/(n+1))
rth arithmetic mean b/w 2x and y: 2x + r*((y-2x)/(n+1))
since both are equal
x + r*((2y-x)/(n+1)) =2x + r*((y-2x)/(n+1))
=> r * (((2y-x)-(y-2x))/(n+1)) = x
=> r*(x+y)/(n+1) = x


Thanks and Regards,
Pratik Tibrewal
askiitians faculty
BTech